Two hydrogen atoms moving at the same speed in opposite directions collide up with zero kinetic energy. Each atom then emits light with a wavelength of 121.6 nm. At what speed were the atoms moving before the collision considering that the mass of a hydrogen atom is equal to that of a proton?
1
Expert's answer
2020-10-19T13:20:37-0400

The kinetic energy, "K=\frac{1}{2}mu^{2}" of hydrogen was released in form of photon.

Energy of a photon, "E=\frac{hc}{\lambda}"

Therefore "\frac{1}{2}mu^{2} =\frac{hc}{\lambda}"

"\frac{1}{2} \times1.67\times10^{-27}\times" "u^{2} = \frac{6.626\times10^{-34}\times3\times10^{8}}{121.6\times10^{-9}}"

"u=4.43\times10^{4}ms^{-1}"